OpenStudy (anonymous):

heather invested 20,000 in an account tha pays 6% simple interest if she invests the moeny for 10 years how much will she have

OpenStudy (anonymous):

20000*.06*10=12000

OpenStudy (anonymous):

interest=prt/100 interset=20000*6*10/100=12000 at the end she will have 20000+12000=32000